What's the result of this? 2i * 7+6i=?
Is it: 14i+6? Why? Shouldn't it be: 7+12i^2 ?
Thanks

Answer by Alan3354(69443)

What's the result of this? 2i * 7+6i=?
Is it: 14i+6? Why? Shouldn't it be: 7+12i^2 ?
=============
2i * 7+6i = 2i*7 + 2i*61
= 14i + 12i^2
= 14i - 12
= -12 + 14i